Sodium acetate (Na CH3COO), (pKa=4.8), is added to water. Determine the pH of a 0.015 M solution of sodium acetate. (Hint CH3OO− is the conjugate baseof acetic acid CH3COOH
Answer #1

Sodium acetate is salt of strong base NaOH and weak acid acetic acid.

Given, concentration of salt (C) = 0.015 M.

pKa = 4.8

Now, For salt of weak acid and strong base

pH = 7 +  \frac{1}{2} ( pKa + logC)

Or, pH = 7 + \frac{1}{2} ( 4.8 + log 0.015)

Or, pH = 7 + \frac{1}{2} ( 4.8 - 1.8)

Or, pH = 7 + 1.5

Or, pH = 8.5